f(x)=3x+2, what is the value if x=4?
Found 2 solutions by Alan3354, b4algebra:
Answer by Alan3354(69443)

f(4)=3x+2
-------------
That makes no sense.
4 is not a variable
Do you mean f(x)=3x+2 ?
Then, what's your question?

Answer by b4algebra(1)

if you meant f(x)=3x+2 then you sub in 4 in to the equation=
f(4)=3(4)+2
f(4)=12+2
f(4) equals then 14